The length of a rectangle is 5cm more than twice the width. The preimeter of the rectangle is 34cm. Find the fimensions of the rectangle.
Answer by tiger(2)

let the width of trhe rectangle = x
so, length = x+5
so, perimeter :- 2{(x+5)+x}=34
2(2x+5)=34
4x+10=34
4x=34-10
x=24/4=6
so length=11
width =6